Heights of the students in a class are given in the distribution below:




5' - 5'2"      10




5'2" - 5'4" 40




5'4" - 5'6"   25 .




Find:




1. Central Tendency






1
Expert's answer
2022-03-16T17:08:32-0400

Solution

Central tendency is the measurement of mean, median and the mode.




(a) Mean.

"\bar X=\dfrac{\sum fX}{\sum f}=\dfrac{400.5}{75}=5.34"


(b) median

"\dfrac{N}{2}=\dfrac{75}{2}=37.5"


Hence the median lies in the class 5.2-5.4

Median"=L+(\dfrac{\frac{N}{2}-pcf}{f})i"


"=5.2+(\dfrac {37.5-11}{40})0
.2"


"=5.3325"

Hence the median height is "5'3""


(c) Mode

The class with the highest frequency (40) is 5.2-5.4 and hence is the modal class.


Mode"=L(\dfrac{\Delta_1}{\Delta_1+\Delta_2})i"


"M_0=5.2(\dfrac{(40-10)}{(40-10)+(40-25)})0.2"


"=5.333"


Modal height is "5'3""